### v1

# QNM families: classification and competition

Abstract
The perturbation spectra of black hole (BH) beyond standard general relativity (GR) may exhibit complex structures with long-lived modes that generate echo-like modulations on the ringdown signal. However, a systematic framework for understanding the internal structure of such spectra, the physical nature of different mode families, and their dynamical excitation is still undeveloped. In this paper, we address this issue by proposing a potential methodology that combines frequency-domain classification with time-domain analysis, using a hairy Schwarzschild BH that admits a double-peak perturbative potential as a theoretical platform. Our analysis of the quasinormal mode (QNM) spectrum identifies two distinct families of modes: the photon sphere (PS) family, arising from delocalized scattering resonances, and the echo family, corresponding to highly localized quasi-bound states. We then develop a windowed energy analysis framework in the time domain, which discloses a dynamic competition for dominance between these families. In particular, our results explicitly show that this competition is sensitive to the properties of the initial perturbation source, and that higher-overtone echo modes can dominate in the observed signal, which are in contrast to the standard PS mode in GR. This study establishes the dynamic evolution of this energy competition as a new observational signature for probing new physics and further motivate a supplemental framework for analyzing long-lived ringdown signals.
I Introduction
The dawn of the gravitational-wave (GW) astronomy era, inaugurated by the landmark observations of the LIGO collaborations LIGOScientific:2016aoc , has opened an unprecedented window into the strong field regime of gravity. Future space-based detectors such as LISA Finn:2000sy , Taiji Hu:2017mde , and TianQin TianQin:2015yph and more promise to expand this window even further. One important scenario in this field is black hole spectroscopy Berti:2009kk ; Konoplya:2011qq ; Berti:2025hly , the idea that by analyzing the “ringdown” signal of a post-merger BH, one can probe its intrinsic properties analogous to how atomic spectroscopy reveals the structure of atoms. Within the standard paradigm, the ringdown waveform is modeled as a superposition of QNMs. These modes, characterized by their complex frequencies, are considered carriers of information about the final BH’s mass and spin, providing a powerful framework to test the foundational no-hair theorem of GR Berti:2005ys ; Dreyer:2003bv . This standard picture, however, becomes significantly more complex when considering that the final compact object may not be a Kerr BH, and its perturbation spectrum could exhibit a much richer structure.
Indeed, recent theoretical explorations have shown that the emergence of new QNM families is a widespread phenomenon in a variety of physical scenarios. These new families can be categorized by their physical origins. Firstly, modifications to spacetime’s asymptotic or topological structure can introduce new QNM families. A prime example is found in Cardoso:2017soq , where, in addition to the familiar photon sphere (PS) family tied to the BH’s light ring, a distinct and long-lived de Sitter family arises from the presence of the cosmological horizon. Exotic compact objects like wormholes Cardoso:2016rao ; Bueno:2017hyj ; Qian:2025lau , or even specific braneworld models Tan:2024qij , can feature effective potential wells that trap perturbations, giving rise to long-lived echo families Cardoso:2019rvt . Additionally, the echo modes were also found and analyzed by imposing a non-smooth effective potential Liu:2021aqh ; Daghigh:2025wcw ; Xie:2025jbr . Secondly, the coupling between gravitational perturbations and other matter fields invariably leads to mixed spectra. For examples, the gravito-electromagnetic coupling in Kerr-Newman spacetimes splits the QNM spectra into PS and near-horizon families Dias:2015wqa ; and the gravitational and scalar modes in various scalar-tensor theories become intertwined Guo:2022rms ; Blazquez-Salcedo:2016enn ; Lin:2025lsz . Lastly, the intrinsic properties of the perturbing fields themselves, such as the mass of a Proca field, can generate quasi-bound states that form their own spectral families Rosa:2011my ; Lei:2023wlt . All these studies suggest that complex spacetime geometries or matter contents could leave their imprints directly on the global structure of the QNM spectra.
Since the perturbation spectrum can exhibit such rich structures, it is significant to take care of the global structure of the spectra besides the QNM spectra themselves. Especially, the spectra structure itself, including the number of mode families, their physical origins, hierarchical arrangement, and interaction, constitutes a new and largely untapped channel of information about the spacetime geometry and the underlying physical theory. Cardoso et al. have laid the groundwork by performing a physical classification of QNM families in the frequency domain Cardoso:2017soq , however, the dynamical picture of how these distinct families are excited and compete for dominance as the time goes by remains an open issue. This is important because the temporal evolution of the spectral structure is very crucial for a complete extraction of information from the ringdown signal. Our study aims to bridge this gap by proposing a preliminary but systematic methodology for analyzing complex systems that possesses multiple QNM families. Encouraged by the above, we select an theoretical testbed, a hairy Schwarzschild black hole constructed via the gravitational decoupling (GD) method Ovalle:2020kpd , which naturally gives rise to a double-peak effective potential capable of supporting both PS and echo spectrum families.
Our methodology consists of a two-step workflow. The first step is a physical classification of modes in the frequency domain, which is made by WKB approximation, and we verify the reliability of the results of the QNM spectra with a set of methods. The second step is a dynamical analysis in time domain. To this end, we compute the complex amplitude of the master scalar Berti:2007dg ; Nollert:1998ys and develop a windowed energy analysis framework. By calculating the energy fraction of each mode as a function of parameters related with a sliding observation window, we construct a dynamic portrait of how different families compete for dominance and reveal the sensitivity of this competition to the initial perturbation source.
This paper is organized as follows. In Sec. II, we derive general master equations of hairy Schwarzschild black hole for perturbation field spins $s=0,1,2$, having the effective potential we scan such black hole parameter space and find the phase diagram to distinct single-peak and double-peak region. In Sec. III, we present the hierarchical structure of the multi-families QNM spectrum in the frequency domain and provide a physical classification based on the QNM eigenfunction spacial configurations. In Sec. IV, we detail our windowed time-domain analysis, showcasing the energy competition between QNM families and its source sensitivity. Finally, we summarize our findings and discuss their implications for future observations in Sec. V. Additionally, we review the GD approach in Appendix A, and demonstrate the supplemented derivation for general master equation in Appendix B. Details of the numerical methods for solving the master equation in the frequency domain and time domain are presented in Appendices CE.
II Perturbation equation and echo parameter space
We start from analyzing the perturbation equations, especially the effective potential to give the echo parameter space of the hairy Schwarzschild black hole Ovalle:2020kpd . The construction of this black hole from GD approach and its extended studies are briefly reviewed in Appendix A. To facilitate our study, we copy here the metric (32) of hairy Schwarzschild black hole
$\mathrm{d}s^{2}=-A(r)\mathrm{d}t^{2}+\frac{\mathrm{d}r^{2}}{B(r)}+C(r)(\mathrm{d}\theta^{2}+\sin^{2}\theta\mathrm{d}\phi^{2})\,$, (1)
$\text{with}\quad A(r)=B(r)=1-\frac{2M}{r}+\alpha\,e^{-\frac{r}{M-l_{0}/2}}\equiv f(r)\quad\text{and}\quad C(r)=r^{2}\,$, (2)
This black hole was found to admit an additional PS in certain parameter regimes Guo:2022ghl , comparing against the single PS in Schwarzschild black hole. These multiple PSs bring in richer and distinctive optical appearances Meng:2023htc and interferometric signatures Wang:2025hzu . The presence of additional PS has also found to closely associated with the echo signals of the probe scalar field Cavalcanti:2022cga ; Yang:2024rms . Moreover, the QNM and greybody factor for various perturbations of this black hole has been studied in Yang:2022ifo , in which the authors focus on the case with single peak in the perturbative potentials.
Here, we shall carefully explore the QNM of perturbation around the hairy Schwarzschild black hole (1), especially, when the perturbative potential has double peaks. The master equation for massless perturbations with different field spins can be incorporated into the Regge-Wheeler-like formula
$\Big(-\frac{\partial^{2}}{\partial t^{2}}+\frac{\partial^{2}}{\partial r_{\star}^{2}}-V_{s}\Big)\Psi_{s,\ell,m}=0\,,\quad\text{with}\quad V_{s}(r)=f(r)\Big[\frac{\ell(\ell+1)}{r^{2}}+(1-s)\frac{f^{\prime}(r)}{r}+\delta_{s,2}f^{\prime\prime}(r)\Big]\,$, (3)
where $\mathrm{d}r_{\star}=\mathrm{d}r/f(r)$ denotes the tortoise coordinate, $\Psi_{s,\ell,m}$ denotes the master field reduced from dynamics variables of a certain perturbation equation. Whereas $s=0,1,2$ indicate scalar, electromagnetic and gravitational axial perturbations, respectively, and $\delta_{s,2}$ stands for the Kronecker delta.
To establish a direct correspondence with the properties of the PS, our subsequent analysis of QNMs will concentrate on electromagnetic perturbations $(s=1)$. Although the dynamics of all massless perturbations are known to converge to that of null geodesics Cardoso:2008bp , as one can see that the master equations (3) with different spins become identical in the eikonal limit $(\ell\gg 1)$, the QNMs of electromagnetic fields often provide a better numerical approximation to the geometric optics results for finite, low-lying multipoles.
Before proceeding, we shall briefly review the reducing procedure of the master scalar $\Psi_{1,\ell,m}$. For $s=1$, the perturbation could be sourced by an external electromagnetic field, of which the dynamics is depicted by the Maxwell equation
$\nabla_{\mu}F^{\mu\nu}=g^{\mu\alpha}\nabla_{\mu}(\partial_{\alpha}A^{\nu}-\partial_{\nu}A_{\alpha})=0\,$, (4)
where the field strength tensor $F^{\mu\nu}$ is defined in terms of the vector potential $A_{\mu}$ as $F=\mathrm{d}A$. In a spherically symmetric background, the vector potential can be decomposed into a Regge-Wheeler-Zerilli(RWZ) form
$A_{\mu}=\sum_{\ell,m}\left[\begin{pmatrix}0\\ 0\\ A_{\ell,m}(t,r)\,\frac{1}{\sin\theta}\,\partial_{\varphi}Y_{\ell m}\\ -A_{\ell,m}(t,r)\,\sin\theta\,\partial_{\theta}Y_{\ell m}\end{pmatrix}_{\text{odd}}+\quad\begin{pmatrix}J_{\ell,m}(t,r)\,Y_{\ell,m}\\ H_{\ell,m}(t,r)\,Y_{\ell,m}\\ K_{\ell,m}(t,r)\,\partial_{\theta}Y_{\ell,m}\\ K_{\ell,m}(t,r)\,\partial_{\varphi}Y_{\ell,m}\end{pmatrix}_{\text{even}}\right]\,$, (5)
where for axial modes with odd parity $(-1)^{\ell+1}$, we can simply set
$A_{\ell,m}=\Psi^{\text{odd}}_{1,\ell,m}\,.$ (6)
For polar modes with even parity $(-1)^{\ell}$, the master scalar can be taken as
$\dot{K}_{\ell,m}$ $=$ $\sqrt{A\,B}\,\dot{\Psi}^{\text{even}}_{1,\ell,m}+J_{\ell,m}=f\,\dot{\Psi}^{\text{even}}_{1,\ell,m}+J_{\ell,m}\,$,
$\dot{H}_{\ell,m}$ $=$ $\sqrt{\frac{A}{B}}\,\frac{\ell(\ell+1)}{C}(\Psi^{\text{even}}_{1,\ell,m})^{\prime}+J_{\ell,m}^{\prime}=\frac{\ell(\ell+1)}{r^{2}}(\Psi^{\text{even}}_{1,\ell,m})^{\prime}+J_{\ell,m}^{\prime}\,$, (7)
where the dot represents the time derivative, and the prime represents the radial derivative. The decomposition (5) divides the Maxwell equation into odd part and even part, and introducing of master scalar (6)-(II) then reduces the both equations into a same master equation (3) with $\Psi^{\text{odd}}_{1,\ell,m}=\Psi^{\text{even}}_{1,\ell,m}=\Psi_{1,\ell,m}$. Noted that the parallel procedures for other spins are shown in Appendix B.
We move on to investigate the morphology of the effective potentials. A crucial aspect of the hairy black hole solution (1) is the rich phenomenology that arises from its full parameter space. While the analysis in Cavalcanti:2022cga ; Avalos:2023jeh focused on a specific subspace by imposing the constraint $l_{0}=2M\alpha/(e^{2}+\alpha)$ which adjusts the horizon radius of the hairy black hole horizon radius to that of Schwarzschild black hole (e.g. $r_{h}(M,\alpha,l_{0})=2\mathsf{M}_{s}$). Here we shall explore the broader implications by treating the hairy charge $l_{0}$ and the deformation intensity $\alpha$} as independent variables. The interplay between these two parameters can induce a non-trivial structure in the spacetime geometry. Specifically, for certain regions in the $\{\alpha,l_{0}\}$ parameter space, the metric function $f(r)$ develops two inflection points outside the horizon. This non-monotonicity of the gravitational potential is directly imprinted onto the perturbative dynamics, causing the effective potential $V_{s}(r)$ to exhibit three local extrema. We then scan the parameter $\{\alpha,l_{0}\}$ and map the morphology of the electromagnetic potential with fixed $s=1$, $\ell=6$ and $M=1$ (same settings for later studies) in Fig. 1, where the grey shaded region highlights the parameter combinations that give rise to the double-peak potential, in contrast to the blank region corresponding to the typical single-peak case.
Figure 1: (a): Parameter spaces of the effective potential morphology for electromagnetic perturbations with $s=1$, $\ell=6$ and $M=1$. The grey shaded region indicates the $\{\alpha,l_{0}\}$ parameter space where the potential exhibits a double-peak structure, a prerequisite for echoes. The blank region corresponds to a typical single-peak potential. The black curve shows the constraint $l_{0}=2M\alpha/(e^{2}+\alpha)$ where the hairy black hole’s horizon coincides with the seed Schwarzschild radius. The horizontal and vertical red dashed lines mark the potential varying along typical parameter paths analyzed in (b) and (c), respectively. (b): Typical potential profiles along the horizontal path from (a) at a fixed $\alpha=7$. This path demonstrates the transition from the Outer-Single-Peak regime (left end) to the Inner-Single-Peak regime (right end) as $\alpha$ increases. (c): Typical potential profiles along the vertical path from (a) at a fixed $l_{0}/M=0.94$. This path demonstrates the transition from the Inner-Single-Peak regime (bottom end) to the Outer-Single-Peak regime (top end) as $l_{0}$ increases.
To further explore the phenomenological consequences of this transition, we must select specific paths through this parameter space for our subsequent QNM analysis. To this end, we choose two representative trajectories, one horizontal with fixed $l_{0}/M=0.94$ and one vertical with fixed $\alpha=7$, that traverse the borders between the single-peak and double-peak regions, as indicated by the red dashed lines in Fig. 1. Furthermore, as we see from Fig. 1 and Fig. 1 that the single peak regions can be further classified into Inner-Single-Peak and Outer-Single-Peak depending on which peak exists as the red dashed lines extend beyond the double-peak region. By analyzing the QNM spectra along these paths, we aim to uncover the distinctive phenomenological signatures associated with the emergence of the potential well, providing a direct link between the spacetime geometry and its observable ringdown properties.
III Photon sphere and echo QNM families
Upon performing a temporal Fourier decomposition on the master scalar, $\Psi_{s,\ell,m}(t,r)=e^{-\mathrm{i}\omega t}\Psi_{s,\ell,m}(r)$, the master equation (3) transforms into a one-dimensional Schrödinger-like equation for the complex frequency $\omega$, namely
$\mathcal{M}[\Psi]\equiv\left[\frac{\mathrm{d}^{2}}{\mathrm{d}r_{\star}^{2}}+\omega^{2}-f(r)\left(\frac{\ell(\ell+1)}{r^{2}}+(1-s)\frac{f^{\prime}(r)}{r}+\delta_{s,2}f^{\prime\prime}(r)\right)\right]\Psi_{s,\ell,m}(r)=0\,.$ (8)
The dissipative nature of the black hole spacetime imposes specific boundary conditions on the radial function $\Psi_{s,\ell,m}(r)$. Classically, a QNM is defined as a solution that is purely ingoing at the event horizon ($r_{\star}\to$-$\infty$) and purely outgoing at spatial infinity ($r_{\star}\to\infty$) Nollert:1999ji ; Berti:2009kk ; Konoplya:2011qq . In our present study, the effective potential $V_{s}$ vanishes both at the event horizon and spatial infinity. This translates to the asymptotic behavior of $\Psi_{s,\ell,m}(r)$, i.e.,
$\Psi_{s,\ell,m}(r_{\star}\rightarrow-\infty)\sim e^{-\mathrm{i}\omega r_{\star}},\quad\text{and}\quad\Psi_{s,\ell,m}(r_{\star}\rightarrow\infty)\sim e^{\mathrm{i}\omega r_{\star}}.$ (9)
Solving for QNMs involves dealing with the boundary conditions and the divergence of eigenfunctions at the boundaries. The hyperboloidal framework offers a powerful approach to overcome these difficulties Zenginoglu:2007jw ; Ansorg:2016ztf ; PanossoMacedo:2018hab ; Zenginoglu:2011jz ; PanossoMacedo:2023qzp ; Zenginoglu:2024bzs ; PanossoMacedo:2024nkw ; Jaramillo:2020tuu . By using constant-$\tau$ slices that extend to null infinity and penetrate the event horizon, the hyperboloidal framework prevents the divergence of eigenfunctions. Moreover, at the boundaries, light cones are oriented outward from the computational domain, so the boundary conditions are naturally satisfied. The divergent boundary conditions (9) are replaced by a requirement of global regularity for the eigenfunction across the compactified hyperboloidal slices.
The QNM spectra we obtain by using the pseudospectral method, which is well-suited to the hyperboloidal framework, reveals a remarkably rich structure, particularly in the parameter space near the echo-generating region. This finding motivates us to carry out a more thorough investigation, employing a range of complementary computational techniques to cross-validate the results and to build a complete physical picture. As we will see soon that the double-peak potential in the hairy Schwarzschild black hole leads to a clear bifurcation of the spectrum. Aided by the WKB approximation as a classification tool, we distinct them as PS family and echo family, respectively, meaning that the deformation source not only shift the QNM spectra but also fundamentally alters the organization and classification of the spectra families. Thus, instead of the individual QNMs and their collection in traditional study, here we shall focus more on the global structure of the perturbation spectrum.
III.1 Methodology for QNM calculation and validation
To obtain a reliable picture of the QNM spectra, we employ a suite of distinct methods, each with its own advantage, allowing for cross-validation. We briefly list them as follow and more details are elaborated in the appendices.
  1. 1.
    Pseudospectral Method (Primary Method): Our results are primarily obtained using the pseudospectral method, which transforms the differential operator eigenvalue problem into a matrix eigenvalue problem with the Chebyshev-Lobatto grid being used. The hyperbolidal framework of this hairy black hole is presented in Appendix C.
  2. 2.
    WKB Approximation (for Classification): The WKB method, pioneered by Schutz and Will Schutz:1985km , provides a semi-analytic tool. Its key idea is to connect the solution near a potential extremum with the asymptotic wave solutions. For a potential barrier with $V^{\prime\prime}(r_{0})<0$ at the peak $r_{0}$, matching solutions across the peak yields the well-known quantization condition for complex QNM spectra
    $\frac{Q_{0}}{\sqrt{2Q_{0}^{\prime\prime}}}=\mathrm{i}\Big(n+\frac{1}{2}\Big)\,$, (10)
    where $Q(r_{\star})\equiv\omega^{2}-V(r_{\star})$, and $Q_{0}$, $Q_{0}^{\prime\prime}$ are evaluated at the peak. Since $Q_{0}^{\prime\prime}=-V_{0}^{\prime\prime}>0$, this formula naturally yields a complex $\omega$. More details can be referred to Konoplya:2003ii , however, we shall give some comments on its application on a potential well with $V^{\prime\prime}(r_{0})>0$ at the well $r_{0}$. Since the term $Q_{0}^{\prime\prime}=-V_{0}^{\prime\prime}$ then becomes negative, the square root $\sqrt{2Q_{0}^{\prime\prime}}$ thus produces an imaginary factor, which cancels the $i$ on the right-hand side, leading to a condition for a purely real frequency. Furthermore, the asymptotic WKB solutions far from the turning points take the form,
    $\Psi_{\text{WKB}}(r_{\star})\approx Q(r_{\star})^{-1/4}\exp\Big(\pm\mathrm{i}\int\sqrt{Q(t)}\mathrm{d}t\Big).$ (11)
    For a potential well, a bound state requires $\omega^{2}<V(r_{\star}\to\pm\infty)$, causing $Q(r_{\star})$ to become negative in the asymptotic regions. Consequently, $\sqrt{Q}$ becomes imaginary, and the oscillatory solution transforms into an exponentially decaying one, characteristic of a bound state. This dual reasoning explains why the WKB approximation interprets modes trapped in a potential well as purely real bound states.
  3. 3.
    Direct Integration Method (Shooting Method): In this approach, the master equation (3) is integrated from the boundaries, and the QNM spectra are determined by requiring a smooth match of the solutions in an intermediate region. It is known to be very robust for solving quasi-bound states (echo modes), and its reliability for computing bound states has also been highlighted recently Volkel:2025lhe . Here we mainly employ this method to validate the imaginary part of the echo modes, thus offering a crucial verification of their damping rates. For a detailed account of our implementation, see Appendix D.
  4. 4.
    Time-Domain Analysis with Matrix Pencil: We perform time-domain evolutions of the master equation and extracted the QNM spectra from the late-time ringdown signal using the Matrix Pencil (MP) method Berti:2007dg ; 56027 ; 370583 . This technique is known for its high accuracy and provides a completely independent validation route. A brief overview is given in Appendix E.
III.2 Analysis of the QNM spectrum and method comparison
Using the methods listed above, we calculate the QNM spectra for selected parameter sets with fixed $\alpha=7$, showing in Tab. 1, which provides a detailed look at the QNM spectra for three representative cases and highlights the agreement between different computational methods. More importantly, it reveals the structural changes in the spectra as the effective potential transiting between single-peak and double-peak configurations.
Overtone (n) Pseudospectral Time-Domain (MP) 9th-WKB
Inner-Single-Peak $l_{0}/M=0.890$ 0 1.444794-0.074881 i 1.444794-0.074881 i -
1 1.554661-0.167495 i 1.554658-0.167496 i -
2 1.679244-0.264020 i 1.679484-0.264193 i -
3 3.289470-0.286500 i 3.289550-0.286691 i 3.289452-0.286485 i
[…]
Outer-Single-Peak $l_{0}/M=1.040$ 0 1.276954-0.079749 i 1.276954-0.079749 i 1.276953-0.079748 i
1 1.270596-0.239538 i 1.270539-0.239520 i 1.270537-0.239525 i
2 1.254843-0.374128 i 1.256677-0.399826 i 1.257770-0.400132 i
3 1.280259-0.570814 i 1.216388-0.571014 i 1.238783-0.562110 i
Table 1: Comparison of electromagnetic QNM spectra $(M\omega)$ for selected parameter sets with $\alpha=7$. The results in the table show the agreement for the QNM spectra between pseudospectral method and various validation techniques. The symbol ‘-’ indicates the WKB formula is blind for those overtones, while a purely real value indicates a bound-state prediction for a potential well.
For the Inner-Single-Peak case $(l_{0}/M=0.890)$, the fundamental mode predicted by the WKB method does not correspond to the fundamental mode $(n=0)$ of the pseudospectral results. Instead, it aligns with a higher overtone $(n=3)$. This noteworthy feature indicates that the presence of a new mode family dominates the low-damping part of the spectrum. This feature is robustly confirmed by the Time-Domain (MP) method, ruling out the possibility that these are spurious modes of the pseudospectral method.
In contrast, for the Outer-Single-Peak case ($l_{0}/M=1.040$), the WKB fundamental mode matches the pseudospectral fundamental mode $(n=0)$. This scenario represents the familiar behavior expected from black holes with a typical potential with a single peak, where the spectrum is dominated by a single family of modes associated with the main potential barrier.
The double-peak case ($l_{0}/M=0.952$) showcases the most interesting phenomenology. These modes, characterized by their extremely low decay rate and a distinct echo period of $\mathcal{T}=\frac{2\pi}{\Delta\text{Re}(\omega_{n})}$, are identified by both the Time-Domain (MP) and pseudospectral methods, with the results showing agreement. Furthermore, we employ the WKB approximation around the potential well, which gives a series of purely real, bound-state-like modes. They show good agreement with the real parts of the extremely long-lived echo modes found by other two methods.
The above analysis hints the existence of a new family of QNMs and highlights the efficacy of the WKB method as a mode classifier. This motivates us to scan the parameter space to specify the classification of the QNM spectrum.
III.3 Classification and evolution of QNM families
We scan the parameters along two representative trajectories that traverse the echo parameter space defined by the red dashed lines in Fig. 1. The results of QNM spectra are presented in Fig. 2, where the top panel corresponds to the vertical traversal with fixed $\alpha=7$ while the bottom panel is for horizontal traversal with fixed $l_{0}/M=0.94$. It is worth emphasizing that the QNM spectra obtained from each method are continuous functions of the black hole parameters. Nevertheless, in Fig. 2, we represent the QNM frequencies with distinct discrete markers, solely to facilitate their visual distinction and classification.
Figure 2: Top Panel: The imaginary (left) and real (right) parts of the QNM spectra as $l_{0}/M$ is varied at fixed $\alpha=7$. This corresponds to a vertical traversal of the parameter space shown in Fig. 1. Bottom Panel: The imaginary (left) and real (right) parts of the QNM spectra as $\alpha$ is varied at a fixed $l_{0}/M=0.94$. This corresponds to a horizontal traversal. In each panel, the vertical red dashed lines mark the boundaries between the single-peak and double-peak (echo) regions. Blue Dots: High-precision data from the pseudospectral method, filtered by a spectral convergence test (drift ratio $\geq 10^{3}$ Chen:2024mon ; BOYD199611 ). Colored Polygons: The squares, triangles, and other polygons represent the first mode estimated by WKB approximation obtained from different extrema of the effective potential. They serve to classify the pseudospectral method results into: the PS family (OPS mode for the Outer-Single-Peak, IPS mode for the Inner-Single-Peak) and the echo family (associated with the potential well). Purple Diamonds: These points are calculated using the direct integration method. They provide an independent validation for the imaginary part of the long-lived echo modes, where the WKB approximation is not applicable. The excellent agreement confirms the accuracy of our pseudospectral method results for the damping rates of these modes.
The WKB approximation serves as a guide to the physical origin of the different mode families. As seen in the top panels of Fig. 2, in the Outer-Single-Peak region, the WKB approximations for the Outer Photon Sphere (OPS) modes (red squares) obtained from the outer peak show excellent agreement with the pseudospectral method results (blue dots). This represents the familiar GR-like scenario. However, upon entering the center of double-peak region, the WKB markers for both the OPS and Inner Photon Sphere (IPS) modes obtained from the inner peak lose their accuracies. This failure is expected, as the WKB approximation breaks down when the potential develops a complex, multi-extremum structure.
An interesting phenomenon occurs during the transition from the double peaks to the Inner-Single-Peak region. Unlike the abrupt disappearance of echo modes when transitioning to the Outer-Single-Peak region, the echo family extends smoothly into the Inner-Single-Peak domain. By a quick glance on Fig. 1, we argue that this is because the disappearance of the outer peak leaves behind a prominent “shoulder” in the effective potential. This shoulder creates a region where the potential decays slowly towards spatial infinity, leading to adiffraction trapping effect for waves. This underlying physical mechanism can be intuitively understand when one focus on the spacial configuration of a single mode, as it can be explicitly seen in Fig. 3, where we present the fundamental eigenfunction ($\Psi_{s,\ell,m}=\Psi_{1,6,0}$, blue curves) , computed by the direction integration method, against their related effective potentials ($V_{s}=V_{1}$, yellow curves) for each case.
Figure 3: Left Panel: In the Inner-Single-Peak region, the spacial configuration of the fundamental mode (with $M\omega_{0}=1.405711-0.033302\penalty 10000\ \mathrm{i}$) is significantly condensed near the potential’s “shoulder”, demonstrating the diffraction trapping effect. Middle Panel: In the double-peak region, the spacial configuration of the fundamental mode (with $M\omega_{0}=1.121228-1.569230\times 10^{-7}\penalty 10000\ \mathrm{i}$) is highly localized within the potential well, exhibiting a clear quasi-bound state (QBS) character. Right Panel: In the Outer-Single-Peak region, the spacial configuration of the fundamental mode (with $M\omega_{0}=1.246082-0.096051\penalty 10000\ \mathrm{i}$) is highly delocalized, which is a characteristic of a standard scattering resonance associated with the PS. In each panel, the black dashed lines indicate the energy level, $\text{Re}(\omega_{n})^{2}$, of the corresponding QNM from the pseudospectral data.
In the Outer-Single-Peak case (right panel of Fig. 3), which resembles the standard GR scenario, the eigenfunction is highly delocalized, spreading over a wide spatial range and its energy level $\text{Re}(\omega_{0})^{2}$ approaches the peak of potential barrier. This is the hallmark of a typical scattering resonance, where the wave interacts briefly with the potential barrier before efficiently dissipating to the boundaries, corresponding to a short-lived mode with a large imaginary part.
In contrast, the double-peak case (middle panel of Fig 3) features a highly localized eigenfunction, whose energy level, $\text{Re}(\omega_{0})^{2}$, lies above the bottom of the potential well. As a result, its amplitude almost entirely confined to this region. This is a clear signature of a quasi-bound state (QBS), where the wave is effectively trapped, leading to a long-lived echo mode with an extremely small imaginary part.
The Inner-Single-Peak case (left panel of Fig. 3) represents a fascinating intermediate state. The eigenfunction, of which the energy level $\text{Re}(\omega_{0})^{2}$ approaches the potential’s “shoulder”, is partially localized, showing a significant condensation of its amplitude near this “shoulder.” This configuration perfectly illustrates the diffraction trapping mechanism, a mode that is more confined than a simple scattering resonance but more dissipative than a typical QBS. Thus, analyzing the eigenfunction’s degree of localization provides a potential method for classifying the QNM families.
This re-classification also explains the reordering of the spectral hierarchy observed in Fig. 2. In the Inner-Single-Peak region, the partially trapped echo family becomes the set of fundamental modes. Consequently, the original IPS modes, being highly delocalized scattering resonances, are relegated to highly damped, excited states. The fate of these IPS modes exhibits a subtle dependence on the path taken through the parameter space. Along the vertical traversal (top panel of Fig. 2), the WKB approximation (green triangles) can still track the IPS modes as they extend continuously from the double-peak region into the Inner-Single-Peak region. However, along the horizontal traversal (bottom panel of Fig. 2), the WKB approximation for the IPS modes vanishes abruptly at the transition boundary. This occurs because as the parameters move deeper into the Inner-Single-Peak region along this path, the damping rate of these IPS modes increases dramatically, making them numerically challenging to resolve and requiring higher grid resolution to maintain their convergences above our threshold.
IV Complex amplitude and energy fraction
The analysis in the previous section revealed a rich structure, particularly the competition between the PS and echo QNM families. Then natural questions arise: when two or more QNM families coexist, which one is the dominant mode and how to determine their dominances? An usual treatment is to consider the fundamental modes with the smallest absolute of imaginary part as the dominant mode, but these modes may only become dominant at very late times, and a waveform reconstructed from it would miss crucial information from the early-time signal. Thus, it was addressed in Giesler:2019uxc that including overtones allows for the modeling of the black hole ringdown signal for all times, stressing the importance of the overtone. On the other hand, realistic observation of long-live ringdown signals may inevitably be “windowed”, as such mode’s characteristic lifetime can be extraordinarily long. For instance, using the fundamental echo mode data from Tab. 1, their lifetimes for a supermassive hairy Schwarzschild black hole with $M\sim 10^{7}M_{\odot}$ would be approximately $\tau=\frac{1}{|\text{Im}(M\omega)|}\frac{GM}{c^{3}}\approx 9.9$ years, which can be comparable to the entire mission duration of space-based detectors.
Thus, it is significant to develop the windowed analysis of the ringdown waveform. To start with, let us briefly review the standard procedure of fitting ringdown waveform in a data-driven framework. We shall firstly solve the master equation (3) using the finite difference method (see Appendix E.1 for details) to obtain the time-domain profile of the master scalar $\Psi(t)$. Second, we model the resulting time-series data based on the linear superposition of damped exponents ansatz, $\Psi(t)\simeq\sum_{k=1}^{p}h_{k}\,e^{-\mathrm{i}\omega_{k}t}$, where the sum over $p$ includes all relevant modes $\{\omega_{n}\}$ and their mirror modes $\{-\overline{\omega_{n}}\}$. Assuming that the modeling ringing phase starts at the peak of the signal $t=t_{\text{peak}}=0$ and ends at $t=N\,\Delta t_{\text{eff}}$, where $\Delta t_{\text{eff}}$ is the sampling time step and $N\geq 2p-1$ is an integer, we can write the discrete time-series signal $\Psi(m\,\Delta t_{\text{eff}})\equiv x[m]$ as
$x[m]=\sum_{k=1}^{p}h_{k}(z_{k})^{m}\,,\quad\text{with}\quad z_{k}=e^{-\mathrm{i}\omega_{k}\Delta t_{\text{eff}}}\,$, (12)
where $h_{k}$ is the complex amplitude of the $k$-th mode for the master scalar $\Psi(t)$. To connect this amplitude to a physical observable, we note that the outgoing EM radiation at infinity is described by the Maxwell-Newman-Penrose scalar $\mathbf{\phi}_{2}$ Teukolsky:1973ha , and as demonstrated by Chandrasekhar Chandrasekhar:1985kt , the RWZ master scalar $\Psi$ and the $\mathbf{\phi}_{2}$ are linked by a well-established transformation involving a first-order differential operator. For any given mode $k$ oscillating as $e^{-\mathrm{i}\omega_{k}t}$, this relationship manifests in the frequency domain as a direct proportionality to the mode’s frequency. Therefore, the complex amplitude $h_{k}$ of the master scalar can be related to the QNM excitation coefficient $C_{k}$ Berti:2025hly . As addressed in Andersson:1996cm ; Leaver:1986gd ; Berti:2006wq ; Oshita:2021iyn , the QNM excitation coefficient $C_{k}=E_{k}T_{k}$ can be further factorized into a source-independent excitation factor $E_{k}$ and a source-dependent term $T_{k}$. The factor $E_{k}$ is determined by the residue of the Green’s function at the corresponding pole, while $T_{k}$ depends on the initial data and on the specific source exciting the perturbations Andersson:1996cm . More studies on the QNM excitation coefficient can also be seen in Oshita:2025ibu ; Oshita:2024wgt ; Lo:2025njp and references therein. Here we focus on the complex amplitude $h_{k}$ for our following purpose. As suggested in Berti:2007dg , one can robustly extract $\{\omega_{k}\}$ via applying the MP method and subsequently determine $\{h_{k}\}$ using a linear least-squares fit, the main steps of which are present in Appendix E.2 and Appendix E.3, respectively.
IV.1 Defining energy fraction in a sliding observation window
Since the observation window is always finite, so beside the complex amplitude $h_{k}$ or excitation coefficient $C_{k}$, the decay of a mode is also important for the received ringdown signal. The latter is significant if the system has long-lived modes, especially for the echo mode family found in the current hairy Schwarzschild black hole. Thus, in order to combine both effects in determining the dominant mode, we shall define the new energy fraction to quantify the contribution of each mode within a finite window.
To proceed, we consider a finite observation window starts at $t_{0}=d\,\Delta t_{\text{eff}}$ and ends at $t_{\text{end}}=t_{0}+T=(d+\mathcal{N})\,\Delta t_{\text{eff}}$, and then propose a time-shifting transformation
$t\rightarrow t^{\prime}$ $=$ $t-t_{0}\,$, (13)
$m\rightarrow m^{\prime}$ $=$ $m-d\,.$ (14)
Thus, such window profile for the shifted time coordinate becomes $t^{\prime}\in[0,T]$. Subsequently, the original signal (12) at indices $m^{\prime}=0,1,2,\cdots,\mathcal{N}$ can be rewritten as
$x[m^{\prime}+d]=\sum_{k=1}^{p}h_{k}(z_{k})^{(m^{\prime}+d)}=\sum_{k=1}^{p}(h_{k}z_{k}^{d})(z_{k})^{m^{\prime}}\equiv x_{\text{obs}}[m^{\prime}]=\sum_{k=1}^{p}\tilde{h}_{k}\,(z_{k})^{m^{\prime}}\,.$ (15)
Here $\tilde{h}_{k}$ is indicated as an effective amplitude measured by the observation signal $x_{\text{obs}}[m^{\prime}]$ at $m^{\prime}=0$, which is the primary observable quantity accessible to us. Especially, when the observation window is large enough, we have the relation
$\tilde{h}_{k}(t_{0})\equiv h_{k}(z_{k})^{d}=h_{k}e^{-\mathrm{i}\omega_{k}t_{0}}\,$, (16)
and its magnitude satisfies $|\tilde{h}_{k}|=|h_{k}|e^{-|\text{Im}(\omega_{k})|t_{0}}$, which encodes the decay of the mode up to the start of our observation. Therefore, the energy radiated by a single QNM in an observation window which is proportional to the time-integral of the square of the field’s time derivative Berti:2006hb , can be reduced as
$E_{k}(t_{0},T)$ $\propto$ $\int_{t_{0}}^{T+t_{0}}\left|\frac{\mathrm{d}\Psi_{k}(t)}{\mathrm{d}t}\right|^{2}\mathrm{d}t=\int_{t_{0}}^{T+t_{0}}|h_{k}(-\mathrm{i}\omega_{k})e^{-\mathrm{i}\omega_{k}t}|^{2}\mathrm{d}t$ (17)
$=$ $\int_{0}^{T}|\tilde{h}_{k}(t_{0})(-\mathrm{i}\omega_{k})e^{-\mathrm{i}\omega_{k}t^{\prime}}|^{2}\mathrm{d}t^{\prime}=|\tilde{h}_{k}(t_{0})|^{2}|\omega_{k}|^{2}\int_{0}^{T}e^{-2|\text{Im}(\omega_{k})|t^{\prime}}\mathrm{d}t^{\prime}$
$=$ $|\tilde{h}_{k}(t_{0})|^{2}|\text{Re}(\omega_{k})|^{2}\left(1+\frac{1}{4Q_{k}^{2}}\right)\int_{0}^{T}e^{-2|\text{Im}(\omega_{k})|t^{\prime}}\mathrm{d}t^{\prime}\approx|\tilde{h}_{k}(t_{0})|^{2}|\text{Re}(\omega_{k})|^{2}\int_{0}^{T}e^{-2|\text{Im}(\omega_{k})|t^{\prime}}\mathrm{d}t^{\prime},$
where in the second line we have used (13) and (16) , while in the third line we have used the exact relation $|\omega_{k}|^{2}=|\text{Re}(\omega_{k})|^{2}(1+1/(4Q_{k}^{2}))$, where $Q_{k}=|\text{Re}(\omega_{k})|/(2|\text{Im}(\omega_{k})|)$ is the quality factor. For the QNM modes relevant to the current study, this quality factor is large. For instance, even for a relatively high overtone such as $n=8$ in our echo data, we find $Q_{k}^{2}\approx 92$ and the echo data from Guo:2022umh is $Q_{k}^{2}\sim 10^{8}$. Thus, the correction term $1/(4Q_{k}^{2})$ is of the order of $10^{-3}$ or even smaller, such that the approximation in the last step can be safely employed. Further performing the final integral in (17) shall give us the energy radiated by a single mode within the window
$E_{k}(t_{0},T)\propto\frac{|\tilde{h}_{k}(t_{0})|^{2}|\text{Re}(\omega_{k})|^{2}}{2|\text{Im}(\omega_{k})|}\Big(1-e^{-2|\text{Im}(\omega_{k})|T}\Big)\,.$ (18)
It is noted that in the idealized limits where the observation starts at the physical onset of the ringing $(t_{0}\to 0)$ and the window is long enough to capture the entire signal $(T\to\infty)$, our expression (18) reduces to the standard formula for the total radiated energy for a single multipole, $E_{\text{rad}}\propto|h_{k}|^{2}|\text{Re}(\omega_{k})|^{2}/|\text{Im}(\omega_{k})|$ Berti:2006hb .
Sbusequently, we define the Windowed Ringing Energy Fraction, $\mathcal{E}_{k}(t_{0},T)$, for each mode as the ratio of its energy to the total energy summed over all contributing modes within the window,
$\mathcal{E}_{k}(t_{0},T)\equiv\frac{E_{k}(t_{0},T)}{\sum_{j}E_{j}(t_{0},T)}=\frac{|\tilde{h}_{k}(t_{0})|^{2}|\text{Re}(\omega_{k})|^{2}|\text{Im}(\omega_{k})|^{-1}(1-e^{-2|\text{Im}(\omega_{k})|T})}{\sum_{j}|\tilde{h}_{j}(t_{0})|^{2}|\text{Re}(\omega_{j})|^{2}|\text{Im}(\omega_{j})|^{-1}(1-e^{-2|\text{Im}(\omega_{j})|T})}\,$, (19)
which measures the normalized contribution of each QNM to the ringdown signal observed within the specific ringing phase duration $T$, allowing for an effective comparison of their excitation strengths under the ringdown simulation.
To close this subsection, we summarize our windowed analysis workflow of the ringdown waveform as
[…]
along which we shall quantify the competition between QNM families within finite time windows in the next subsection.
IV.2 Competition for dominant modes
To provide a comprehensive characterization of the competition between the echo and PS QNM families as the time goes by, we move on to analyze three distinct scenarios corresponding to different typical profiles of the double-peak potential.
  1. 1.
    Outer-Peak Dominant Case: The profile of effective potential in this scenario corresponds to the second panel of Fig. 1 as well as the inset of Fig. 4(a) in tortoise coordinates. A typical time domain waveform is shown in Fig. 4(a) with $\alpha=6.78$ and $\ell_{0}/M=0.940$, based on which we depict the energy fraction in Fig. 4. Also, the QNMs and coefficients extracted at $t_{0}=0$ are listed in Tab. 2, where $n=3$ mode corresponds to the WKB-identified OPS mode. By tuning $t_{0}$, we see that the PS mode (red solid curve in Fig. 4) is the most dominant component of the signal. Only at very late time, say around $t_{0}\approx 123M$, the fundamental echo mode ($n=0$) finally gain dominance. This demonstrates a clear temporal competition between the families. Noted that this competition cannot be realized from the frequency-domain analysis in Fig. 2, where with the same parameters, the echo family always occupies the low-lying states.
    To evaluate the influence of the perturbation source on this competition, we also investigate the system’s response to a wider initial Gaussian wave packet, as shown in Fig. 4. Comparing to Fig. 4, the competition still exists as expected, only the dominance of the PS family last longer, saying approximately $134M$. This finding is significant, as it implies that information about the perturbation source is encoded not only in the excitation of individual QNMs but also in the hierarchical structure and competitive dynamics of the entire perturbation spectrum.
    (a) Time-domain waveform
    Figure 4: Time-domain analysis for the Outer-Peak dominant case: (a) shows the time-domain waveform, in which the inset shows the outer-peak dominant potential, the initial Gaussian wave packet, and the observer’s position (vertical dashed line). The main plot’s vertical red dashed line indicates the position of the observation window at $t_{0}=0$, with a fixed length of $T=280M$. (b) and (c) show the windowed energy fraction $\mathcal{E}_{k}$ for each mode as a function of the window’s start time $t_{0}$. The red and blue dashed lines indicate $\mathcal{E}_{k}=1$ and $\mathcal{E}_{k}=0.8$, respectively. The latter marks the onset of the dominant mode. Curves are truncated when the relative error between the extracted frequency and pseudospectral method result exceeds 0.7%. (b) is sourced from the initial Gaussian wave $\Psi(t=0,r_{\star})=\exp[-(r_{\star}-1075)^{2}/8]$ following (a), while (c) is sourced from a wider wave packet: $\Psi(t=0,r_{\star})=\exp[-(r_{{\star}}-1075)^{2}]/72]$.
    n QNMs (MP) Error Complex Coefficients
[…]
    6 $\pm$1.369064 - 0.140374 i $4\times 10^{-4}$ 0.182142 $\pm$ 0.910646 i
[…]
    Table 2: QNMs and complex coefficients $\{\tilde{h}_{k}\}$ extracted at $t_{0}=0$ in Fig. 4(a). The error is the relative difference with the pseudospectral method result. The $n=3$ mode (highlighted in red in Fig. 4) corresponds to the WKB-identified PS mode.
  2. 2.
    Inner-Peak Dominant Case: The profile of effective potential in this scenario corresponds to the last second panel of Fig. 1 and the inset of 5 in tortoise coordinates. The typical results in this case are shown in Fig. 5 and Tab. 3 for $\alpha=7.30$ and $l_{0}/M=0.940$.
    Fig. 5 shows that the PS mode (purple curve, $n=4$) is excited with a very low initial energy fraction. Consequently, as the analysis window slides, this mode’s contribution diminishes rapidly, leading to its anomalous disappearance from the detectable signal even before the higher-overtone echo mode ($n=5$). In contrast, when the system is perturbed by the narrower packet in Fig. 5, the early energy fraction of the PS mode is significantly boosted. As a result, it persists longer in the signal and decays in the expected hierarchical order, after the $n=5$ mode.
    This suggests that, in the present scenario, not only is the competition among different QNM families highly sensitive to the specific perturbation source, but the individual family also exhibit different sensitivities to distinct perturbation source, since it alters the initial energy distributed into higher-frequency modes, resulting an observable effect at late time. Our findings suggest that future ringdown simulations could potentially leverage the differential response of competing QNM families to extract valuable information about the perturbative sources.
    Figure 5: Time-domain analysis for the Inner-Peak dominant case with $\alpha=7.30$ and $l_{0}/M=0.940$. Other settings are similar to those in Fig. 4, but here (c) is sourced from a narrower wave packet: $\Psi(t=0,r_{\star})=\exp[-9(r_{\star}-1075)^{2}/8]$.
    n QNMs (MP) Error Complex Coefficients
[…]
    2 $\pm$1.538542 - 0.114471 i $\times 10^{-6} 11$9.949436 $\pm$ 391.803857 i
[…]
    Table 3: QNMs and complex coefficients $\{\tilde{h}_{k}\}$ extracted at $t_{0}=0$ in Fig. 5. The error is the relative difference with the pseudospectral method result. The $n=4$ mode corresponds to the WKB-identified PS mode.
  3. 3.
    Equal-Height Peaks Case: In this scenario, the two potential peaks are of comparable height (see the fourth panel of 1 and the inset of 6), which represents the most trapping mechanism and the perturbation spectrum is almost dominated by the echo family. This allows us to study the dynamic competition within the echo family itself. The results are shown in Fig. 6 and Tab. 4 with $\alpha=6.88$ and $l_{0}/M=0.940$. Comparing to the previous cases, the echo signal in Fig. 6 is explicit.
    As illustrated in Fig. 6, the energy fraction reveals an obvious competition among the different overtones of echo QNM family. At very late time ($t_{0}\approx 112M$), the $n=2$ mode (green curve) surprisingly emerges as the dominant component of the signal. Although it eventually cedes its dominance to the lower overtones, for a finite observation window, an observer would perceive this $n=2$ excited state as the dominant mode. These findings mean that in ringdown analysis, even at what might be considered as late time, higher-overtone modes can still be dominant, due to the fact that the late-time portion of the signal within the window does not necessarily correspond to the late-time stage of the entire ringing process. Thus, when dealing with long-lived modes, one must be particularly cautious in defining what constitutes the late time in the rindown process, especially, in context of long-live-mode detection, one cannot simply fit the “late-time signal within the observation window” by sorely applying the fundamental mode.
    This picture becomes even more intricate when we consider a wider initial wave packet, see Fig. 6. It shows that the $n=2$ mode can still achieve dominance near a local peak where $t_{0}\approx 106M$, but then begin yielding to the $n=1$ mode . The comparison in this case highlights the conclusion that the late-time behaviors of long-lived modes are highly sensitive to the initial properties of the perturbation source, which is in contrast to the shorter-lived PS modes generated by a single-peak potential. For PS modes, the late-time ringdown signal is generally considered to be insensitive to the initial configuration and is dominated by the fundamental mode Berti:2005ys ; LIGOScientific:2016lio ; Gossan:2011ha . So, we argue that the different responses of the echo overtones to the source could serve as a potential new channel for extracting information about the astrophysical origin of the perturbation from future gravitational wave observations.
    Figure 6: Time-domain analysis for the Equal-Height Peaks case with $\alpha=6.88$ and $l_{0}/M=0.940$. Other settings are similar to those in Fig. 4(a).
    n QNMs (MP) Error Complex Coefficients
[…]
    Table 4: QNMs and complex coefficients $\{\tilde{h}_{k}\}$ extracted at $t_{0}=0$ in Fig. 6. The error is the relative difference with the pseudospectral result. All the modes are contributed solely from the echo family.
V Conclusions and discussion
In this work, we performed a comprehensive investigation into the eletromagnetic perturbation dynamics of hairy Schwarzschild black holes, which admit both single and double photon spheres. Instead of merely identifying phenomena such as echoes, we establish a physical classification of the QNM spectrum and reveal a new layer of dynamics in the time-domain ringdown signal. Our analysis indicates a rich global structure of QNM spectra arising from the complex structure of the effective potential for this black hole. Our main findings are summarized as follows.
  • With the use of suite of parallel methods, we identified that the QNM spectrum in this scenario generates two physical families: a photon sphere family and an echo family. We proposed that the WKB approximation can serve as an effective marker for these families. Moreover, we argue that the degree of localization of the QNM eigenfunction in the vicinity of the potential barriers can be an unambiguous criterion for their classification, from which we can physically distinguish among the highly delocalized scattering resonances (the PS family), highly localized quasi-bound states (the Echo family), and intermediate diffraction-trapped states observed near border of sing peak and double peaks. In particular, it is seen that from the frequency domain, the strong trapping regime, where the two potential peaks are of comparable height, the traditional PS family predicted by geometric optics is effectively suppressed or absent from the spectra.
  • Motivated by the presence of extremely long-lived modes and in this case an observational data may be inherently windowed, we developed a windowed energy analysis framework to quantitatively study the ringdown signal, which reveals a dynamic competition for dominance among different QNM families. Our results show that the dominant mode within a given analysis window, even one starting at what is considered a “late” time, is not necessarily the fundamental mode of the entire spectra. It could instead be a mode from the PS family or even a high-overtone mode from the echo family, depending on the window’s start time.
    Moreover, our studies also indicate that the competitive dynamic for dominance is profoundly sensitive to the properties of the initial perturbation source. In particular, the initial energy distribution imparted by the source dictates the hierarchy and evolution of the competing modes. This feature is different from the short-live signal in GR, in which the black hole’s late-time ringdown is independent of the initial excitation LIGOScientific:2016lio ; Gossan:2011ha .
Our current studies can be extended into the gravitational perturbations to analyze the gravitational ringdown signal. We shall argue that our preliminary results can provide significant insights on the observations and analysis of ringdown signal. Firstly, our results indicate that in some cases, higher overtone may become dominant at late time within a certain observational windows. It further confirms that traditional consideration that the observed signal decays into the fundamental mode, might be complemented by an alternative strategy Giesler:2019uxc ; Oshita:2021iyn . Especially, for the system producing long-live signals, developing tailored fitting models, such as those based on a single, pure excited state, is motivated for a more faithful representation of the waveform in the window. Secondly, we suggest that the time-dependent energy competition among QNM families may itself serve as a distinctive observable. In particular, the temporal re-ordering of dominant modes during ringdown window could provide potentially robust way of differing hairy black holes from their counterpart in GR, helping to further test no-hair theorem with ringdown overtone signal Bhagwat:2019dtm . The last but not least, the spectral evolution of the ringdown, including the dynamic distribution and underlying transfer of energy among QNM families, encodes rich information of both the spacetime geometry (the profile of perturbation potential) and mechanism driving the perturbation. This raises the possibility of using the ringdown’s temporal structure to probe more information about the wave sources.
Additionally, it is known that in the eikonal limit of GR and many other theories, $\text{Re}(\omega)$ corresponds to the orbital frequency of an unstable null geodesic at the PS, while $\text{Im}(\omega)$ corresponds to its Lyapunov exponent, which quantifies the characteristic timescale of the orbit’s instability Cardoso:2008bp ; Dolan:2010wr . However, here in the case of a double-peak potential, particularly when the peaks are of similar height, the QNM spectrum is almost governed by the echo family. This corresponds to the eigenfunction configurations in the middle panel of Fig. 3. The energy levels of echo modes $\text{Re}(\omega_{n})^{2}$ are high enough to reach the potential’s peak, which possesses an energy of an unstable orbit on the PS in the geometric-optics picture. Nevertheless, their decay rates given by $\text{Im}(\omega_{n})$ are determined by their overtone numbers $n$ within the echo family, which is different from the decay rate predicted by the WKB approximation at the local maximum of the potential, related to the Lyapunov exponent. In a word, since for black holes with double PS, the QNMs corresponding to the PS are either absent or extremely suppressed, therefore, the link inferring the geometric-optics properties like the Lyapunov exponent from the QNM spectra breakdowns. It suggests that instead of the geometric-optics approximation, the QNM structure of these black holes could require a more rigorous framework, which can fully incorporates the wave effects of trapping and resonant tunneling. This deserves further efforts. It is worthwhile to note that the authors of Duran-Cabaces:2025sly recently attempted to infer the Lyapunov exponent from the properties of extra rings formed by double-peak potentials, but found significant discrepancies with QNM calculations. Our current study could offer a parallel physical explanation for this discrepancy.
Acknowledgement
This work was supported in part by the National Natural Science Foundation of China with Grants Nos. 12375054 and 12505067, and the Postgraduate Research & Practice Innovation Program of Jiangsu Province with Grant No. KYCX25_3923.
Appendix A Hairy Schwarzschild black hole from gravitational decoupling
In this appendix, we shall briefly review how the hairy Schwarzschild black hole can be constructed by the GD algorithm, which was developed for analytically finding a deformed solution to Einstein’s field equation with additional source term Ovalle:2018gic ; Ovalle:2017wqi ; Ovalle:2020kpd . The methodology can be started from taking the additional source as a decoupled subpart of a total energy-momentum tensor $\tilde{T}_{\mu\nu}$, and the Einstein equations take the form
$\tilde{G}_{\mu\nu}\equiv R_{\mu\nu}-\frac{1}{2}Rg_{\mu\nu}=\kappa\,\tilde{T}_{\mu\nu}\equiv\kappa(T_{\mu\nu}+\Theta_{\mu\nu})\,$, (21)
where $T_{\mu\nu}$ represents the energy-momentum tensor corresponding to a known seed GR solution, while $\Theta_{\mu\nu}$ describes the contributions from new matter fields or additional gravitational sectors. The simplest seed solution is the Schwarzschild vacuum, given by $T_{\mu\nu}=0$, then one can consider a deformed Schwarzschild solution to (21) with the ansatz
$\mathrm{d}s^{2}=-\Big(1-\frac{2\mathsf{M_{s}}}{r}\Big)e^{\alpha\,k(r)}\mathrm{d}t^{2}+\left(1-\frac{2\mathsf{M_{s}}}{r}\right)^{-1}e^{-\alpha\,k(r)}\mathrm{d}r^{2}+r^{2}\mathrm{d}\Omega^{2}\,$, (22)
where $\mathsf{M_{s}}$ represents the Schwarzschild mass and $k(r)$ is a function to be determined. The parameter $\alpha$ quantifies the deformation strength. $\alpha=0$ indicates no additional source $\Theta_{\mu\nu}$, reverting the solution to the original seed metric without deformation. It is worth noting that this ansatz becomes trivial deformation when the function $k(r)$ is satisfied with
$\quad e^{\alpha\,k(r)}=1-\frac{\alpha\,c_{0}}{r}\left(1-\frac{2\mathsf{M_{s}}}{r}\right)^{-1}\penalty 10000\ \penalty 10000\ \text{with}\penalty 10000\ \penalty 10000\ c_{0}=\text{Const.}\,$, (23)
because it leads to
$-g_{tt}=g_{rr}^{-1}=1-\frac{2\mathsf{M_{s}}+\alpha\,c_{0}}{r}=1-\frac{2\mathsf{M^{\prime}_{s}}}{r}\,$, (24)
which leaves the original Schwarzschild geometry unaffected, and only the mass parameter will be changed.
[…]
$\tilde{p}_{r}=-\tilde{\rho}\quad\text{or}\quad\Theta^{r}{}_{r}=\Theta^{t}{}_{t}\,$, (25)
where the effective density, radial pressure and tangential pressure can be taken from
$\tilde{T}^{\mu}{}_{\nu}\equiv\text{Diag}\{-\tilde{\rho},\tilde{p}_{r},\tilde{p}_{t},\tilde{p}_{t}\}=T^{\mu}{}_{\nu}+\Theta^{\mu}{}_{\nu}\,$. (26)
Noted that in the current framework we already have $T^{\mu}{}_{\nu}=0$. Secondly, the strong energy condition is imposed to avoid exotic matter sources, which guarantees the convergence of time-like geodesics affected by the tidal force of matter sources, requiring
$\tilde{\rho}+\tilde{p}_{r}+2\tilde{p}_{t}\geq 0\,,\penalty 10000\ \penalty 10000\ \tilde{\rho}+\tilde{p}_{r}\geq 0\,,\penalty 10000\ \penalty 10000\ \tilde{\rho}+\tilde{p}_{t}\geq 0\,.$ (27)
Subsequently, the “quasi-Einstein” equations combining with the above conditions yield two second-order linear differential inequalities
$\frac{1}{2\kappa\,r}\Big[h^{\prime\prime}(r-2\mathsf{M_{s}})+2h^{\prime}\Big]$ $=$ $\Theta^{\theta}{}_{\theta}\geq 0\,$, (28)
$\frac{1}{2\kappa\,r^{2}}\Big[h^{\prime\prime}(r-2\mathsf{M_{s}})r+4h^{\prime}\mathsf{M_{s}}-2h+2\Big]$ $=$ $\Theta^{\theta}{}_{\theta}-\Theta^{t}{}_{t}\geq 0\,$, (29)
where $h\equiv e^{\alpha\,k(r)}$ and the prime denotes the derivative with respect to $r$. One may consider the deformation should be trivial at $r=r_{h}$ and $r\gg\mathsf{M_{s}}$. Since the deformed black hole admits a proper horizon $r_{h}\sim 2\mathsf{M_{s}}$ and a asymptotically Schwarzschild geometric, thus a simple example emerges as
$\Theta^{\theta}{}_{\theta}=\frac{\alpha}{2\kappa\,\mathsf{M_{s}}^{2}r}e^{-r/\mathsf{M_{s}}}\left(r-2\mathsf{M_{s}}\right)\,.$ (30)
Therefore, the deformation function sourced by (30) can be solved from the second-order differential equation (28)
$h(r)=c_{1}-\alpha\frac{c_{2}-r\,e^{-r/\mathsf{M_{s}}}}{r-2\mathsf{M_{s}}}=1-\left(\frac{l_{0}}{r}-\alpha\,e^{-\frac{r}{\mathsf{M_{s}}}}\right)\left(1-\frac{2\mathsf{M_{s}}}{r}\right)^{-1},$ (31)
where the integration constant $c_{1}=1$ has been taken for keeping asymptotical flatness and $\alpha\,c_{2}\equiv l_{0}$ has been defined for representing a hairy charge associated with primary hair.
Subsequently, a hairy black hole solution deformed from the Schwarzschild seed metric reads Ovalle:2020kpd
$\mathrm{d}s^{2}=-A(r)\mathrm{d}t^{2}+\frac{\mathrm{d}r^{2}}{B(r)}+C(r)(\mathrm{d}\theta^{2}+\sin^{2}\theta\mathrm{d}\phi^{2})\,$,
$\text{with}\quad A(r)=B(r)=1-\frac{2M}{r}+\alpha\,e^{-\frac{r}{M-l_{0}/2}}\equiv f(r)\quad\text{and}\quad C(r)=r^{2}\,$, (32)
where $M=\mathsf{M_{s}}+l_{0}/2$ represents the asymptotic mass of the hairy black hole in this framework. It is noted that, in general, different choice of $\Theta_{\mu\nu}$ could result to different hairy black hole holes, but the hairy Schwarzschild black hole (1) was obtained by considering $\Theta_{\mu\nu}$ as the stress energy tensor of an anisotropic fluid. Nevertheless, this hairy black hole still has its generality because it does not rely on specific matter fields and accommodates a wide variety of hair, such as scalar hair, tensor hair, and even fluid-like dark matter, etc.. Thus, some theoretical aspects and observational phenomena of this hairy black hole and its rotating counterpart have been widely studied. For example, the thermodynamical properties were studied in Mahapatra:2022xea . The observational signatures through strong gravitational lensing, black hole shadow and images were explored in Afrin:2021imp ; Islam:2021dyk ; Ghosh:2022kit ; Afrin:2021wlj ; Meng:2024puu ; Meng:2025ivb which are based on the deflection of photons. Starting from the geodesic motion of timelike particle, the effects such as precession and Lense-Thirring effect Wu:2023wld , gravitational waves from extreme mass ratio inspirals Zi:2023omh and quasi-periodic oscillations Liu:2023ggz have also been examized.
Appendix B Reducing procedure of the master scalar with other spins
In this Appendix, we will briefly shown the reducing procedure of the master scalar (3) for the scalar field ($s=0$) and gravitational perturbation ($s=2$).
  • For $s=0$, the perturbation could be sourced by an external scalar field and then be governed by the Klein-Gorden equation in the black hole background
    $\square\Phi=\frac{1}{\sqrt{-g}}(g^{\mu\nu}\sqrt{-g}\Phi_{,\mu})_{,\nu}=0\,$, (33)
    in which $g$ denotes the determinant of the background. Considering a standard separation of variables of the form
    $\Phi(t,r,\theta,\phi)=\sum_{\ell,m}\frac{\Psi_{0,\ell,m}(t,r)}{r}Y_{\ell,m}(\theta,\phi)\,$, (34)
    it is straightforward to obtain that the master scalar $\Psi_{0,\ell,m}$ will involve via the master equation (3) with $s=0$, and the standard scalar spherical harmonics $Y_{\ell,m}$ is governed by
    $\left[\frac{1}{\sin\theta}\frac{\partial}{\partial\theta}\left(\sin\theta\frac{\partial}{\partial\theta}\right)+\frac{1}{\sin^{2}\theta}\frac{\partial^{2}}{\partial\phi^{2}}\right]Y_{\ell,m}(\theta,\phi)=-\ell(\ell+1)Y_{\ell,m}(\theta,\phi)\,.$ (35)
  • For $s=2$, we consider a simple case that perturbations are excited by a particle falling radially into the black hole. Due to the spherical symmetry, the resulting scalar perturbations are of pure even-parity. Consequently, the odd-parity gravitational perturbations naturally decouple and evolve independently at the linear level. Their dynamics are thus governed by a homogeneous Einstein equation, whose solutions reveal the intrinsic dynamic of the spacetime itself
    $G^{(1)}_{\mu\nu}[h^{\text{odd}}]$ $=$ $\left(\delta_{\mu}{}^{\alpha}\delta_{\nu}{}^{\beta}-\frac{1}{2}g_{\mu\nu}g^{\alpha\beta}\right)R^{(1)}_{\alpha\beta}[h^{\text{odd}}]=0\,$, (36)
    $\text{with}\quad h^{\text{odd}}_{\mu\nu}$ $=$ $\begin{pmatrix}0&0&-h_{0}(t,r)\,\csc\theta\,\partial_{\phi}&h_{0}(t,r)\,\sin\theta\,\partial_{\phi}\\ 0&0&-h_{1}(t,r)\,\csc\theta\,\partial_{\phi}&h_{1}(t,r)\,\sin\theta\,\partial_{\phi}\\ -h_{0}(t,r)\,\csc\theta\,\partial_{\phi}&-h_{1}(t,r)\,\csc\theta\,\partial_{\phi}&0&0\\ h_{0}(t,r)\,\sin\theta\,\partial_{\phi}&h_{1}(t,r)\,\sin\theta\,\partial_{\phi}&0&0\end{pmatrix}Y_{\ell,m}\,$, (37)
    where $G^{(1)}_{\mu\nu}$ and $R^{(1)}_{\alpha\beta}$ denote the linear Einstein tensor and linear Ricci tensor. The odd-parity metric perturbation, $h^{\text{odd}}_{\mu\nu}$, is constructed from the Regge-Wheeler basis functions and can be expressed in terms of two radial functions, $h_{0}(t,r)$ and $h_{1}(t,r)$. Following the standard procedure, we introduce a master scalar function $\Psi^{\text{odd}}_{2,\ell,m}$ to encapsulate their dynamics. A convenient definition relates the master scalar to the metric functions as follows:
    $h_{1}$ $=$ $\sqrt{\frac{C}{A\,B}}\,\Psi^{\text{odd}}_{2,\ell,m}=\frac{r}{f}\Psi^{\text{odd}}_{2,\ell,m}\,$, (38)
    $\dot{h}_{0}$ $=$ $\sqrt{A\,B}\,(\sqrt{C}\,\Psi^{\text{odd}}_{2,\ell,m})^{\prime}=f\,(r\,\Psi^{\text{odd}}_{2,\ell,m})^{\prime}\,$, (39)
    the dynamical components $G^{(1)}_{31}[\Psi^{\text{odd}}_{2,\ell,m}]=G^{(1)}_{13}[\Psi^{\text{odd}}_{2,\ell,m}]$ of linear Einstein’s equation (36) can be finally reduced to the master equation (3) with $s=2$.
Appendix C Hyperboloidal framework of hairy Schwarzschild black hole
In this appendix, we employ the so-called “scri-fixing technique” to construct the hyperboloidal coordinate Zenginoglu:2007jw ; PanossoMacedo:2023qzp . Specifically, the hyperboloidal coordinate transformations are given by
$t=r_{h}(\tau-H(\sigma))\,,\quad r=\frac{r_{h}}{\sigma}\,$, (40)
where the length scale has been chosen by the event horizon $r_{h}$. There only exists one horizon for the double-peak effective potential, and the horizon will be $\sigma=1$. From (32), in terms of $\sigma$, the metric function can be expressed as
$\mathcal{F}(\sigma)=f(r(\sigma))=1-\frac{2M}{r_{h}}\sigma+\alpha\exp\Big(-\frac{r_{h}}{(M-l_{0}/2)\sigma}\Big)\,.$ (41)
The derivative of the dimensionless tortoise coordinate $x(\sigma)$ is expanded as
$x^{\prime}(\sigma)=-\frac{1}{\sigma^{2}\mathcal{F}(\sigma)}=-\frac{1}{\sigma^{2}}-\frac{2M}{r_{h}}\frac{1}{\sigma}+\mathcal{O}(1)\,.$ (42)
The leading terms contribute with the singular quantities, i.e.,
$x_{0}(\sigma)=\frac{1}{\sigma}-\frac{2M}{r_{h}}\ln\sigma\,.$ (43)
At the event horizon $\sigma=1$, the integration around the horizon $\sigma=1$ yields
$x_{h}(\sigma)=\frac{1}{K_{h}(1)}\ln|\sigma-1|\,,\quad\text{with}\quad K_{h}(1)=\frac{2M}{r_{h}}+\frac{2\alpha r_{h}}{l_{0}-2M}\exp\Big(\frac{2r_{h}}{l_{0}-2M}\Big)\,.$ (44)
The regular piece $x_{\text{reg}}(\sigma)$ is satisfied with
$x^{\prime}_{\text{reg}}(\sigma)=x^{\prime}(\sigma)-x^{\prime}_{h}(\sigma)-x^{\prime}_{0}(\sigma)\,.$ (45)
Therefore, derivatives of two height functions $H^{\text{in-out}}(\sigma)$ and $H^{\text{out-in}}(\sigma)$ associated with $\sigma$ are
$\frac{\mathrm{d}H^{\text{in-out}}(\sigma)}{\mathrm{d}\sigma}=-\frac{1}{\sigma^{2}\mathcal{F}(\sigma)}+\frac{2}{\sigma^{2}}+\frac{4M}{r_{h}}\frac{1}{\sigma}\,$, (46)
and
$\frac{\mathrm{d}H^{\text{out-in}}(\sigma)}{\mathrm{d}\sigma}=\frac{1}{K_{h}(1)}\frac{2}{\sigma-1}+\frac{1}{\sigma^{2}\mathcal{F}(\sigma)}\,.$ (47)
Note that when $\alpha=0$, i.e., the case of the Schwarzschild black hole, two height functions are equal. Ref. PanossoMacedo:2023qzp tells us how to use the approach presented in it to obtain the hyperboloidal coordinate in a single chart. For some black holes, such as Schwarzschild black hole, RN black hole Destounis:2021lum , quantum corrected Schwarzschild black hole Cao:2024oud , Hayward black hole Wu:2024ldo , and Boulware-Deser-Wheeler black hole Cao:2024sot , a single hyperboloidal chart can indeed be achieved. But things turned out contrary to one’s wishes. We find that within some certain parameters, applying the method provided in PanossoMacedo:2023qzp to this hairy black hole will result in the failure of the construction of hyperboloidal framework using the height function Eq. (46) and Eq. (47). Similar case will happen in the black hole with halo PanossoMacedo:2023qzp . Hence, we need to give some complements for the hyperboloidal framework.
It arises from demanding $g_{\sigma\sigma}>0$, so that the correct signature of metric can be still maintained. The condition $g_{\sigma\sigma}>0$ yields that $|\gamma|<1$, which means that the hypersurfaced $\tau=\text{constant}$ keep spacelike in the concerned domain $\sigma\in(0,1)$ from a geometric perspective. The function $\gamma(\sigma)$ is given by
$\gamma(\sigma)=\frac{\mathrm{d}H(\sigma)}{\mathrm{d}\sigma}p(\sigma)\,.$ (48)
However, after our test, it is found that both strategies derived from PanossoMacedo:2023qzp will encounter situations where condition $|\gamma|<1$ is not met. As some typical examples, in the Fig. 7, we choose some parameters to show that $|\gamma|>1$ appears in the part of the interval $(0,1)$. Although Ref. PanossoMacedo:2023qzp suggests that the out-in strategy is better, what we want to say here is that the out-in strategy is even worse within such chosen of parameters. The situation in the left panel of the Fig. 7 can still be remedied, while the situation in the right panel of the Fig. 7 is already regarded as hopeless. One key difference between these two is the sign of the derivative of function $\gamma(\sigma)$ at the endpoint of the interval including $\sigma=1$. We have $\mathrm{d}\gamma^{\text{in-out}}(\sigma)/\mathrm{d}\sigma(\sigma=1)<0$ but $\mathrm{d}\gamma^{\text{out-in}}(\sigma)/\mathrm{d}\sigma(\sigma=1)>0$ in these examples. Therefore, one can find some interval $(\sigma_{0},1)$, in which the function $\gamma^{\text{in-out}}(\sigma)$ is satisfied with $|\gamma|<1$ in such interval. However, it is not achievable for the function $\gamma^{\text{out-in}}(\sigma)$ at this time.
Now, we provide remedies if a situation similar to the left panel of Fig. 7 occurs. Motivated by Zenginoglu:2007jw , hyperboloidal foliations can be constructed that are spacelike in the interior without affecting the QNM spectra. To accomplish this objective, it is easy to construct the new function $\tilde{\gamma}(\sigma)$ (yellow line in left panel of Fig. 7), which is satisfied with $|\tilde{\gamma}(\sigma)|<1$, where the asymptotic behaviors of $\gamma^{\text{in-out}}(\sigma)$ and $\tilde{\gamma}(\sigma)$ are the same. In practice, if we encounter the above problems, we will adopt $\tilde{\gamma}(\sigma)$ instead of $\gamma(\sigma)$.
Figure 7: The functions $\gamma^{\text{in-out}}(\sigma)$, $\tilde{\gamma}(\sigma)$ and $\gamma^{\text{out-in}}(\sigma)$ are shown. For the left panel with the in-out strategy, the parameters are chosen as $M=1$, $l_{0}=1.02M$ and $\alpha=9.312$. For the right panel with the out-in strategy, the parameters are chosen as $M=1$, $l_{0}=1.02M$ and $\alpha=7.312$. The above parameter selections are from Yang:2024rms .
Using the obtained hyperboloidal coordinates (40), we recast Eq. (3) into a form of a hyperbolic equation
$\partial_{\tau}u=\mathrm{i}Lu\,,\quad L=\frac{1}{\mathrm{i}}\begin{pmatrix}0&1\\ L_{1}&L_{2}\end{pmatrix}\,,\quad u=\begin{pmatrix}\Psi\\ \Pi\end{pmatrix}\,$, (49)
where $\Pi=\partial_{\tau}\Psi$ is introduced to reduce the time derivative order in the equation, and the explicit expressions of differential operators $L_{1}$ and $L_{2}$ are
$L_{1}=\frac{1}{w(\sigma)}\Big[\partial_{\sigma}(p(\sigma)\partial_{\sigma})-q_{s}(\sigma)\Big]\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ L_{2}=\frac{1}{w(\sigma)}\Big[2\gamma(\sigma)\partial_{\sigma}+\partial_{\sigma}\gamma(\sigma)\Big]\,$, (50)
where the explicit forms of the above function are given as follows
$p(\sigma)=\sigma^{2}\mathcal{F}(\sigma)\,,\quad w(\sigma)=\frac{1-\gamma^{2}(\sigma)}{p(\sigma)}\,,\quad q_{s}(\sigma)=\frac{r_{h}^{2}}{p(\sigma)}V_{s}(\sigma)\,.$ (51)
Having the above first-order-in-time PDE system (49), and performing the Fourier transformation in $\tau$ with $u(\tau,\sigma)\sim u(\sigma)e^{\mathrm{i}\omega\tau}$, we arrive at the eigenvalue problem $Lu=\mathrm{i}\omega u$. To numerically compute the QNM spectra, we discretize the differential operator $L$ using a Chebyshev-Lobatto collocation scheme (pseudospectral method), which yields its matrix representation $\mathbf{L}$. This discretization effectively reduces the original infinite-dimensional spectral problem to a finite-dimensional algebraic eigenvalue problem
$\mathbf{L}\mathbf{u}=\mathrm{i}\omega\mathbf{u}\,.$ (52)
Note that we have $\omega_{\tau}=r_{h}\omega_{t}$, in which $\omega_{\tau}$ is derived in the $\tau$-coordinate and $\omega_{t}$ is derived in the $t$-coordinate [cf. Eq. (40)]. Note that when the master equation has a correction term of the velocity, the corresponding hyperbolidal framework can still work Cao:2025qws . Furthermore, the hyperboloidal framework for the Kerr spacetime is constructed in PanossoMacedo:2019npm , which is used to obtain the QNM spectra Cai:2025irl ; Assaad:2025nbv . See also some comments on the hyperboloidal framework for black hole QNMs Qian:2025jht .
Appendix D Direct integration method
In this appendix, we detail the semi-analytic method used to compute the QNM spectra in the frequency domain. The approach is based on finding series solutions to the master equation in the asymptotic regions near the event horizon and spatial infinity. Then matching these solutions in an intermediate region. The fundamental mathematical idea of this direct integration method was originally established by Chandrasekhar Chandrasekhar:1975 . Subsequently, this technique has been widely applied in various contexts, such as in the study of dynamical Chern-Simons gravity Wagle:2021tam ; Molina:2010fb , and the black hole with a thin shell Laeuger:2025zgb .
We begin with Eq. (8). In the spirit of the Frobenius method, we seek solutions in the asymptotic regions by factoring out the singular behavior, which is determined by the QNM boundary conditions (purely ingoing at the horizon and purely outgoing at infinity). At the event horizon $(r\to r_{h})$, the solution takes the form:
$\Psi_{\text{h}}(r)=(r-r_{h})^{-\frac{\mathrm{i}\,\omega}{2\kappa}}\sum^{\mathcal{H}}_{j=0}\mathfrak{h}_{j}(r-r_{h})^{j}\,$, (53)
and at spatial infinity $(r\to\infty)$, the solution is written as:
$\Psi_{\text{inf}}(r)=e^{\mathrm{i}\omega\,r_{\star}(r)}\sum^{\mathcal{I}}_{k=0}\mathfrak{p}_{k}\,r^{-k}.$ (54)
Here, $\{\mathfrak{h}_{j}\}$ and $\{\mathfrak{p}_{k}\}$ are the expansion coefficients at the horizon and infinity, respectively, which are typically functions of the black hole parameters and the frequency $\omega$. $\kappa=f^{\prime}(r_{h})/2$ is the surface gravity at the event horizon, $\mathcal{H}$ and $\mathcal{I}$ are the truncation orders.
To determine these coefficients $\{\mathfrak{h}_{j}\}$ and $\{\mathfrak{p}_{k}\}$, we substitute the series solutions, $\Psi_{\text{h}}(r)$ and $\Psi_{\text{inf}}(r)$ into Eq. (8), i.e., $\mathcal{M}[\Psi_{\text{h}}(r)]$, $\mathcal{M}[\Psi_{\text{inf}}(r)]$. We then expand the resulting expressions around $r_{h}$ and $\infty$. By collecting terms of the same order, we transform the differential equation into two algebraic power series equations:
$\sum^{\mathcal{H}}_{j=0}\mathcal{C}_{j}(\mathfrak{h}_{0},\mathfrak{h}_{1},\cdots,\mathfrak{h}_{j};\omega)(r-r_{h})^{j}=0\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \sum^{\mathcal{I}}_{k=0}\mathcal{B}_{k}(\mathfrak{p}_{0},\mathfrak{p}_{1},\cdots,\mathfrak{p}_{k};\omega)\,r^{-k}=0\,.$ (55)
For these series equations to be valid for all values of $r$ in their respective regions of convergence, the coefficient of each power term must vanish independently. These conditions, i.e., $\mathcal{C}_{j}=0$ and $\mathcal{B}_{k}=0$ for all $j$ and $k$, allow us to derive the following recurrence relations for the expansion coefficients:
$\mathfrak{h}_{j}$ $=\sum^{j-1}_{i=0}\Lambda_{ji}\,\mathfrak{h}_{i}\quad\text{for}\quad j=1,2,\cdots,\mathcal{H}\,$, (56)
$\mathfrak{p}_{k}$ $=\sum^{k-1}_{i=0}\Upsilon_{ki}\,\mathfrak{p}_{i}\quad\text{for}\quad k=1,2,\cdots,\mathcal{I}\,.$ (57)
The coefficients $\Lambda_{ji}$ and $\Upsilon_{ki}$ depend on the black hole parameters, namely $\Lambda_{ji}=\Lambda_{ji}[f^{(n)}(r_{h}),r_{h},\ell,s,\omega,\kappa]$ and $\Upsilon_{ki}=\Upsilon_{ki}[f^{(n)}(\infty),\ell,s,\omega]$. These relations show that each coefficient can be determined from the preceding ones. These two systems of linear equations can be intuitively expressed in the matrix form, where the coefficients form a lower triangular matrix:
$\begin{pmatrix}\mathfrak{h}_{1}\\ \mathfrak{h}_{2}\\ \mathfrak{h}_{3}\\ \vdots\\ \mathfrak{h}_{\mathcal{H}}\end{pmatrix}$ $=\begin{pmatrix}\Lambda_{10}&0&0&0&\cdots&0\\ \Lambda_{20}&\Lambda_{21}&0&0&\cdots&0\\ \Lambda_{30}&\Lambda_{31}&\Lambda_{32}&0&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ \Lambda_{\mathcal{H}0}&\Lambda_{\mathcal{H}1}&\Lambda_{\mathcal{H}2}&\Lambda_{\mathcal{H}3}&\cdots&\Lambda_{\mathcal{H}\,\mathcal{H}-1}\end{pmatrix}\begin{pmatrix}\mathfrak{h}_{0}\\ \mathfrak{h}_{1}\\ \mathfrak{h}_{2}\\ \vdots\\ \mathfrak{h}_{\mathcal{H}-1}\end{pmatrix}\,$, (58)
$\begin{pmatrix}\mathfrak{p}_{1}\\ \mathfrak{p}_{2}\\ \mathfrak{p}_{3}\\ \vdots\\ \mathfrak{p}_{\mathcal{I}}\end{pmatrix}$ $=\begin{pmatrix}\Upsilon_{10}&0&0&0&\cdots&0\\ \Upsilon_{20}&\Upsilon_{21}&0&0&\cdots&0\\ \Upsilon_{30}&\Upsilon_{31}&\Upsilon_{32}&0&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ \Upsilon_{\mathcal{I}0}&\Upsilon_{\mathcal{I}1}&\Upsilon_{\mathcal{I}2}&\Upsilon_{\mathcal{I}3}&\cdots&\Upsilon_{\mathcal{I}\,\mathcal{I}-1}\end{pmatrix}\begin{pmatrix}\mathfrak{p}_{0}\\ \mathfrak{p}_{1}\\ \mathfrak{p}_{2}\\ \vdots\\ \mathfrak{p}_{\mathcal{I}-1}\end{pmatrix}\,$. (59)
Since the expressions for the higher-order coefficients $\Lambda_{ji}$ are very lengthy, we only show the first few $\Lambda_{ji}$ and $\Upsilon_{ki}$ here:
[…]
in which we denote that $f_{r_{h}}^{(n)}=\partial^{(n)}_{r}f|_{r\rightarrow r_{h}}$ and $f_{\infty}^{(n)}=\partial^{(n)}_{r}f|_{r\rightarrow\infty}$. By solving the recurrence relations (56) and (57) iteratively, we can express all higher-order expansion coefficients in terms of the leading-order coefficients, $\mathfrak{h}_{0}$ and $\mathfrak{p}_{0}$:
$\mathfrak{h}_{j}$ $=\mathfrak{h}_{0}\,\mathsf{G}_{j}(\Lambda_{10},\cdots,\Lambda_{j\,j-1})\quad\text{for}\quad j=1,2,\cdots,\mathcal{H}\,$, (61)
$\mathfrak{p}_{k}$ $=\mathfrak{p}_{0}\,\mathsf{G}_{k}(\Upsilon_{10},\cdots,\Upsilon_{k\,k-1})\quad\text{for}\quad k=1,2,\cdots,\mathcal{I}.$ (62)
Due to the structural similarity of the linear systems (58) and (59), the general solutions for the coefficients at the horizon and at infinity share the same functional form, denoted by $\{\mathsf{G}_{i}\}$. However, the truncation orders, $\mathcal{H}$ and $\mathcal{I}$, can be differ. For clarity, we list the explicit forms of the first few functions $\{\mathsf{G}_{i}\}$:
[…]
where $\mathcal{A}_{ij}$ represents either $\Lambda_{ij}$ (for the horizon expansion) or $\Upsilon_{ij}$ (for the infinity expansion).
Substituting these solutions back into the asymptotic expansions and setting $\mathfrak{h}_{0}=\mathfrak{p}_{0}=1$ without loss of generality, we obtain well-defined asymptotic forms of the radial function. Given $\omega\in\mathbb{C}$, the series $\Psi_{\text{h}}(\tilde{r}_{h},\omega)$ and $\Psi_{\text{inf}}(\tilde{r}_{\infty},\omega)$, provide the initial conditions for solving Eq. (8). Here, $\tilde{r}_{h}=(1+\epsilon)r_{h}$ is a point slightly away from the horizon, with $\epsilon$ being a small parameter to avoid the singularity, while $\tilde{r}_{\infty}\gg r_{h}$ is a sufficiently large, which numerically represents infinity. In practice, for a given initial guess for frequency $\omega_{g}$, Eq. (8) can be solved numerically as an initial value problem, starting from both boundaries:
$\mathcal{M}[\Psi_{1}]$ $=0\,,\quad\text{with initial conditions at }\tilde{r}_{h}:\quad\Psi_{1}=\Psi_{\text{h}}(\tilde{r}_{h},\omega_{g})\,,\quad\Psi^{\prime}_{1}=\Psi^{\prime}_{\text{h}}(\tilde{r}_{h},\omega_{g})\,$, (64)
$\mathcal{M}[\Psi_{2}]$ $=0\,,\quad\text{with initial conditions at }\tilde{r}_{\infty}:\quad\Psi_{2}=\Psi_{\text{inf}}(\tilde{r}_{\infty},\omega_{g})\,,\quad\Psi^{\prime}_{2}=\Psi^{\prime}_{\text{inf}}(\tilde{r}_{\infty},\omega_{g})\,.$ (65)
This procedure yields two numerical solutions, $\Psi_{1}(r,\omega_{g})$ and $\Psi_{2}(r,\omega_{g})$, which are obtained by integrating outwards from the horizon and inwards from infinity, respectively. A true QNM spectum of Eq. (8) is found only when these two solutions match smoothly. This matching condition is enforced by requiring their Wronskian to vanish at an arbitrary matching point $r_{m}$ (where $\tilde{r}_{h}<r_{m}<\tilde{r}_{\infty}$):
$\mathcal{W}[\Psi_{1},\Psi_{2}]=\Psi_{1}(r_{m},\omega_{g})\Psi^{\prime}_{2}(r_{m},\omega_{g})-\Psi^{\prime}_{1}(r_{m},\omega_{g})\Psi_{2}(r_{m},\omega_{g})=0\,.$ (66)
Therefore, we find the QNM spectra by treating the Wronskian as a function of the complex frequency $\omega_{g}$ and searching for its roots. We employ a numerical root-finding algorithm, such as Newton’s method, to iteratively refine the guess $\omega_{g}$ until the Wronskian is zero to a desired precision. Once an eigenvalue $\omega_{\text{QNM}}$ is found, the corresponding global eigenfunction can be constructed by joining the two numerical solutions at the matching point, i.e.,
$\Psi(r,\omega_{\text{QNM}})=\begin{cases}\mathcal{N}_{1}\Psi_{1}(r,\omega_{\text{QNM}}),&r\leq r_{m},\\ \mathcal{N}_{2}\Psi_{2}(r,\omega_{\text{QNM}}),&r>r_{m},\end{cases}$ (67)
where $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ are normalization constants chosen to ensure continuity of the function at $r_{m}$. Note that in order to mitigate the influence of finite boundary effects on numerical results, we thus adopt higher truncation orders. In our numerical calculations for the GD hairy black hole, we use the truncation orders $\mathcal{H}=13$ and $\mathcal{I}=14$, with the integration boundaries set at $\epsilon=10^{-4}$ and $\tilde{r}_{\infty}=90M$.
Appendix E Methodology for ringdown analysis
In this appendix, we detail the numerical methodology used to generate the time-domain waveform and subsequently extract the QNM spectra and their corresponding excitation properties. Our approach is a multi-step process designed to first produce a reliable time-series signal and then robustly identify the mode frequencies and their physical contributions.
E.1 Time-domain evolution: finite difference method
To obtain the time-domain profile of the master scalar, we solve the master equation (3) using the finite difference method in time profile. For more details on this standard approach, readers may refer to Zhu:2014sya ; Yang:2023lcm ; Yang:2024cjf .
We begin by discretizing $r_{\star}=j\Delta r_{\star}$ and time coordinate $t=i\,\Delta t$. The master scalar and the potential are evaluated at discrete grid points $\Psi(t,r_{\star})=\Psi(i\,\Delta t,j\,\Delta r_{\star})=\Psi_{i,j}$ and $V(r)=V(r(r_{\star}))=V(r(j\Delta r_{\star}))=V(r_{j})=V_{j}$. The second-order time and space derivatives in the master equation are then approximated using central difference stencils. This transforms the partial differential equation into an iterative scheme that evolves the waveform in time:
$\Psi_{i+1,j}=\frac{\Delta t^{2}(\Psi_{i,j-1}-2\Psi_{i,j}+\Psi_{i,j+1})}{r_{\star}^{2}}-\Delta t^{2}V_{j}\Psi_{i,j}+2\Psi_{i,j}-\Psi_{i-1,j}\,.$ (68)
To ensure the stability of this numerical scheme, the Courant-Friedrichs-Lewy (CFL) condition must be satisfied. For this wave-like equation, the condition requires $\Delta t/\Delta r_{\star}\leq 1$ Smith1985 . We choose our grid spacings to be $\Delta r_{\star}=0.004$ and $\Delta t=0.002$ to safely meet this criterion.
The evolution is initiated with a Gaussian wave packet of the form $\Psi(t=0,r_{\star})=c\cdot\exp[-(r_{\star}-a)^{2}/(2b^{2})]$, with the field being zero for $t<0$. The resulting waveform typically exhibits three distinct phases: an initial outburst from the source, the characteristic damped oscillations of the ringdown phase, and a late-time power-law tail. It is from the ringdown phase that we extract the QNMs and their complex amplitudes using methods described below.
E.2 QNM spectra extraction: matrix pencil method
To determine the signal poles $\{z_{k}\}$, we employ the Matrix Pencil (MP) method, a robust technique known for its high accuracy and stability in the presence of noise Berti:2007dg ; 56027 ; 370583 . This method circumvents the potential instabilities of traditional Prony methods by recasting the problem in terms of a generalized eigenvalue problem.
The procedure begins with constructing two $(N-L)\times L$ Hankel matrices, $Y_{0}$ and $Y_{1}$, from the $N$ data points of the time series $x[n]$. $L$ is the pencil parameter, chosen to be in the range $p\leq L\leq N-p$. The matrices are defined as:
$Y_{0}=\begin{pmatrix}x[0]&x[1]&\cdots&x[L-1]\\ x[1]&x[2]&\cdots&x[L]\\ \vdots&\vdots&\ddots&\vdots\\ x[N-L-1]&x[N-L]&\cdots&x[N-2]\end{pmatrix}$, (69)
and
$Y_{1}=\begin{pmatrix}x[1]&x[2]&\cdots&x[L]\\ x[2]&x[3]&\cdots&x[L+1]\\ \vdots&\vdots&\ddots&\vdots\\ x[N-L]&x[N-L+1]&\cdots&x[N-1]\end{pmatrix}$. (70)
These matrices are then used to form a matrix pencil, $Y_{1}-zY_{0}$. The signal poles $\{z_{k}\}$ are the generalized eigenvalues of the matrix pair $(Y_{1},Y_{0})$ that reduce the rank of the pencil. Numerically, this is achieved by first performing a Singular Value Decomposition (SVD) via $Y_{0}=USV^{T}$ and then estimating the largest singular values of $S$ as our mode order $p$ . The poles $\{z_{k}\}$ are then found as the standard eigenvalues of a reduced $p\times p$ matrix $Z_{p}=S^{-1}_{p}U^{T}_{p}Y_{0}V_{p}$ constructed from the SVD components of $Y_{0}$ and the Hankel matrix $Y_{1}$. Our QNM data is extracted by setting pencil parameter $L=N/2$.
E.3 Complex amplitude determination
With the set of poles $\{z_{k}\}$ determined, we find the complex amplitudes $\{h_{k}\}$ by solving the linear system defined by Eq. (12) using all $N$ available data points. This forms an overdetermined system of linear equations:
$\begin{pmatrix}z_{1}^{0}&z_{2}^{0}&\cdots&z_{p}^{0}\\ z_{1}^{1}&z_{2}^{1}&\cdots&z_{p}^{1}\\ \vdots&\vdots&\ddots&\vdots\\ z_{1}^{N-1}&z_{2}^{N-1}&\cdots&z_{p}^{N-1}\end{pmatrix}\begin{pmatrix}h_{1}\\ h_{2}\\ \vdots\\ h_{p}\end{pmatrix}=\begin{pmatrix}x[0]\\ x[1]\\ \vdots\\ x[N-1]\end{pmatrix}$. (71)
This system, in the form $\mathbf{V}\mathbf{h}=\mathbf{x}$, where $\mathbf{V}$ is a Vandermonde matrix, is solved for $\mathbf{h}$ using a linear least-squares algorithm. This ensures that the solution for $\{h_{k}\}$ is optimal for the entire signal segment under consideration.